# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ G.3.3 Amplification of IGB with depth
$\frac{\textrm{Var}_{\chi}\left({h^{(L)}}\right)}{\textrm{Var}_{\mathcal{W}}\left({\left\langle{h^{(L)}}\right\rangle}\right)}=\mathcal{O}\left(\frac{1}{L}\right)\Longrightarrow\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O^{(c)}}\right)}=\mathcal{O}\left(L\right)\Longrightarrow\lim_{L\rightarrow\infty}\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O^{(c)}}\right)}=\infty\,.$ (210)
We observe a fundamental distinction compared to the previous cases. Network depth can amplify the level of IGB in systems where it is already present. However, depth alone does not induce IGB. Considering a model with a linear activation function and no max pooling after the first hidden layer, and using Gaussian data centered at zero, we have $\left\langle{\rho^{(1;m)}}\right\rangle=\left\langle{h^{(1)}_{m}{}}\right\rangle=0$. Thus, there is no Node Symmetry Breaking (NSB) in the second layer; in other words, the nodes in the second layer are Gaussian r.v.s centered at zero. It can be easily shown that, by iterating this process through each layer, the same scenario will be repeated in each subsequent layer up to the output layer. Therefore, NSB does not occur even in the output nodes, resulting in the absence of IGB in the system.
@@ Appendix H Extension to multi-class problems
We want now to extend the analysis to the multi-class case, i.e. to problems with $N_{C}>2$. Following the framework we introduced earlier, we can easily extend the computation. We will again consider the distribution among the $N_{C}$ classes of the dataset elements after initialization. In particular, we can define $N_{C}$ values associated with the fraction of dataset elements classified as belonging to each of the $N_{C}$ classes, $\left\{f_{i}\right\}_{i=0}^{N_{C}-1}$. We could also consider the distribution of the sorted frequencies; in other words, in each experiment, we order the classes according to the corresponding frequencies and not by their label. To distinguish these frequencies (and their statistics) from the set $\left\{f_{i}\right\}_{i=0}^{N_{C}-1}$ , we use indices $\tilde{i}$, and call this new set $\left\{f_{\tilde{i}}\right\}_{\tilde{i}=0}^{N_{C}-1}$. In particular $f_{\tilde{0}}$ indicates the biggest frequency among the $N_{C}$, $f_{\tilde{1}}$ the second one and so on. Finally we add the set (or subset) cardinality, $M$, on the output nodes set $\left\{O^{(c)}_{M}\right\}_{c=0}^{M-1}$. [^19] Here, the cardinality of the set indicates the number of classes over which the statistics are computed. The statistics of the r.v. change with the number of classes (or nodes), i.e., with the cardinality. Analogously to the set of fractions, we define the set of ranked output nodes $\left\{O^{(\tilde{c})}_{M}\right\}_{\tilde{c}=0}^{M-1}$ In the following, similarly to Sec. E.1 we will derive the distribution of $f_{0}$.
Figure H.1: The distribution $p_{f_{0}}\left({x}\right)$ in a multi-class problem with $N_{C}=10$. The empirical mean value is reported; note that despite the difference in the distribution the mean value is the same that we would observe using a linear activation function $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{N_{C}}$. For these simulations we used the dataset GB on the model SHLP with max pooling ($m=20$). (see App. I.4 for more details).
We will make in the derivation of $f_{0}$ the following approximation: We will repeat the same approach of the binary setting comparing the generic class ”0” with the class with bigger mean output value among the $N_{C}-1$ remaining classes.
The first ingredient that we need is therefore the statistics of $O^{(\tilde{0})}_{N_{C}-1}$. From the analysis of previous sections, we know that $p_{O^{(\tilde{0})}_{N_{C}-1}}^{(\chi)}\left({x}\right)$ is asymptotically Gaussian. The mean value of this Gaussian variable will follow the distribution of the maximum among $N_{C}-1$ drawn from the Gaussian distribution $p_{\mu_{c}}\left({x}\right)$ as, by definition,
$\left\langle{O^{(\tilde{0})}_{N_{C}-1}}\right\rangle\equiv\max_{c\in\{1,\dots,N_{C}\}}\left\langle{O^{(c)}}\right\rangle\,.$ (211)
@@
$p_{Y}\left({y}\right)\equiv p_{\left\langle{O^{(\tilde{0})}_{N_{C}-1}}\right\rangle}\left({y}\right)=\left(N_{C}-1\right)\mathcal{N}\left(y;0,\tilde{\sigma}^{2}\right)\left(\frac{1}{2}\left(1+\operatorname{erf}\left(\frac{y}{\sqrt{2}\tilde{\sigma}}\right)\right)\right)^{N_{C}-2}\,.$ (214)
@@
$p_{f_{0}}\left({x}\right)=\int_{-\infty}^{\infty}p_{\left\langle{O^{(0)}_{N_{C}}}\right\rangle}\left({\tilde{x}}\right)p_{\left\langle{O^{(\tilde{0})}_{N_{C}-1}}\right\rangle}\left({\tilde{x}-\hat{\Delta}\left(x\right)}\right)d\tilde{x}\,$, (215)
where $\left\langle{O^{(\tilde{0})}_{N_{C}-1}}\right\rangle$ the biggest mean output among the $N_{C}-1$. Note that $\hat{\Delta}\left(f_{0}\right)$ is a function of $f_{0}$. Substituting Eq. (212) and Eq. (214) we have:
$p_{f_{0}}\left({x}\right)=\int_{-\infty}^{\infty}\mathcal{N}\left(\tilde{x};0,\hat{\sigma^{2}}\right)(N_{C}-1)\mathcal{N}\left(\tilde{x};\hat{\Delta}\left(x\right),\hat{\sigma^{2}}\right)\left(\frac{1}{2}\left(1+\operatorname{erf}\left(\frac{\tilde{x}-\hat{\Delta}\left(x\right)}{\sqrt{2}\hat{\sigma}}\right)\right)\right)^{N_{C}-2}d\tilde{x}\,$. (216)
@@ H.1 Increasing number of classes exacerbates IGB
Having a large number of classes increases the probability that the center of the distribution related to one of these classes is an outlier of the distribution $p_{\mu_{c}}(x)$. However, since $p_{\mu_{c}}(x)$ is Gaussian (a fast-decaying distribution), the typical value of the maximum increases slowly. Therefore, an extremely large number of classes, $N_{c}$, will be needed to observe significant changes.
We can analyze this quantitatively. The typical value of the maximum out of $N_{c}$ Gaussian random variables of variance $\epsilon$ grows as $\sqrt{(2\epsilon\log(N_{c}))}$ [Gumbel, 1958, Hartarsky et al., 2019]. In the absence of IGB, $\epsilon$ is the variance of $p_{\mu_{c}}(x)$ and scales inversely with the dataset size, i.e., $\sim 1/D$ (see e.g., App. D.2 and Fig. D.1). This implies that outliers will only appear if the number of classes far exceeds the number of datapoints, a situation that does not occur in real life.
@@ I.1.1 A prototype for highly correlated data: MNIST
Figure I.1: Comparison of the distribution $p_{f_{0}}\left({x}\right)$ between two untrained neural networks fed with MNIST dataset; the ten classes of the dataset were merged in two macro groups: even and odd. The two architectures employed in the comparison differ in the choice of activation function and pooling. For these simulations we used the E&O dataset on the model CNN-B. (see App. I.4 for more details). On the right, the same plot in a logarithmic scale to better visualize the differences in the low-density regions.
@@ I.1.2 A further example: CIFAR10
Figure I.2: Comparison of the distribution $p_{f_{0}}\left({x}\right)$ between two untrained neural networks fed with CIFAR10 dataset (with 10 classes). The two architectures employed in the comparison differ in the choice of activation function and pooling. In the absence of IGB, the distributions would concentrate around the vertical line. For these simulations we used the dataset C10 on the model CNN-B described below (see App. I.4). On the right, the same plot in a logarithmic scale to better visualize the differences in the low-density regions.
We now show a multi-class example, on the same network as in App. I.1.1 (CNN-B) we propagated CIFAR10. The results are reported in Fig. I.2. As with MNIST, the results are qualitatively consistent with the conclusions of our theory: the network with ReLU and max pooling displays a stronger IGB than its counterpart with $\tanh$ and AvgPooling. Note, also, how in both cases, symmetry between classes is preserved at the ensemble level. In both cases, in fact, we get $\mkern 1.5mu\overline{\mkern-1.5muf_{c}\mkern-1.5mu}\mkern 1.5mu=1/N_{C}$ . To show this we show in Fig. I.2 two vertical lines at the mean values calculated on the empirical distributions (histograms), each surrounded by its own uncertainty, estimated through the standard error.
@@ I.1.3 High cardinality dataset: Cifar100
We also provide another example of a multi-class dataset. In this case, we choose Cifar-100 as a prototype of a high-cardinality dataset (with a high number of classes) to demonstrate the presence of IGB in this scenario as well. In the absence of IGB, the distribution should be peaked at $f_{0}=1/N_{C}=1/100$. However, the histogram shows a peak at 0 (most of the time, the generic class does not receive any assigned elements). The gap between the peak at 0 and the rest of the distribution, particularly evident in the log-log scale plot, indicates a resolution limit due to the finite number of dataset elements ($D=10^{4}$).
Figure I.3: CNN-A (ReLU + max pool case) with C100 in this case ($N_{C}=100$). The peak at $f_{0}=\frac{1}{N_{C}}$, corresponding to the No IGB case (in the absence of IGB, the distribution should concentrate around this value), is reported as a reference. The plot on the right contains the same data as that on the left, but has a logarithmic $x$ axis. We added a small numer, $\epsilon=1e-9$, to the value of $f_{0}$ in order to show the peak at $f_{0}=0$.
Figure I.4: Right: Comparison of the trend of $\max_{i}\{f_{i}\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). The curves show a consistent pattern with the diagram on the shown in Fig. 2. The simulations were conducted on a ResNet (top) and Vision Transformer (bottom) using a binary dataset (dogs vs cats from CIFAR) as input; more details on the setting are provided in App. I.4.
Figure I.5: Right: Comparison of the trend of $\max_{i}\{f_{i}\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). The curves show a consistent pattern with the diagram on the shown in Fig. 2. The simulations were conducted on different architectures (MLP- mixer (top), ResNet (middle) and Vision Transformer (bottom)) using a multi-class dataset (C10) as input; more details on the setting are provided in App. I.4.
@@ I.1.4 Additional architectures
We also provide empirical evidence of IGB in different architectures. To improve the legibility of the paper, we show results for ResNet, ViT and MLP-mix in Sec. I.2. While App. I.2 is devoted to the dynamics und IGB, the reader can recognize the presence of IGB in Figs. I.4 and I.5. Specifically, IGB is identified by checking the quantity $\max_{c}\{f_{c}(\mathcal{W}_{t})\}$, for $t=0$.; we can observe that at time $t=0$, i.e., at initialization, each of these curves shows $\max_{c}\{f_{c}\}\neq 1/N_{C}$. [^21] The reader might have particular interest in the $K=0$ case, where the data is centered and IGB cannot be attributed to data preprocessing.
@@ I.2 Experiments on other architectures & effects on the training dynamics
Even in the case of $K=0$, the system exhibits IGB, as indicated by $\max_{i}\{f_{i}\}\neq 1/N_{C}$ at initialization. Also, it is worth noticing that the susceptibility of networks to IGB varies, as seen from the comparison of the $\gamma(\mathcal{A},\psi(\chi))$ range in the colormaps of different experiments. Despite using the same set of $K$ values, the resulting IGB level varies significantly across networks. The IGB values, dependent on the combined effect of $\psi(\chi)$ and $\mathcal{A}$, show substantial variation among the considered architectures, suggesting the presence of regulatory or amplifying elements for IGB within these architectures.
@@ I.3 IGB in Pre-Trained Models
Figure I.6: The distribution $p_{f_{0}}\left({x}\right)$ on pre-trained models. From left to right are shown PT-ResNet, PT-SwinT-FC, PT-ENetV2-FC on a binary dataset (Cats and Dogs from C10).
@@
Figure I.7: The distribution $p_{f_{0}}\left({x}\right)$ on pre-trained models. From left to right are shown PT-SwinT-MLP, PT-ENetV2-MLP on a binary dataset (Cats and Dogs from C10)

### v5

@@ G.3.3 Amplification of IGB with depth
$\frac{\textrm{Var}_{\chi}\left({Z^{(L)}}\right)}{\textrm{Var}_{\mathcal{W}}\left({\left\langle{Z^{(L)}}\right\rangle}\right)}=\mathcal{O}\left(\frac{1}{L}\right)\Longrightarrow\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O_{c}}\right)}=\mathcal{O}\left(L\right)\Longrightarrow\lim_{L\rightarrow\infty}\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O_{c}}\right)}=\infty\,.$ (210)
We observe a fundamental distinction compared to the previous cases. Network depth can amplify the level of IGB in systems where it is already present. However, depth alone does not induce IGB. Considering a model with a linear activation function and no max pooling after the first hidden layer, and using Gaussian data centered at zero, we have $\left\langle{U^{(1;1)}}\right\rangle=\left\langle{Z^{(1)}{}}\right\rangle=0$. Thus, there is no Node Symmetry Breaking (NSB) in the second layer; in other words, the nodes in the second layer are Gaussian r.v.s centered at zero. It can be easily shown that, by iterating this process through each layer, the same scenario will be repeated in each subsequent layer up to the output layer. Therefore, NSB does not occur even in the output nodes, resulting in the absence of IGB in the system.
@@ Appendix H Extension to multi-class problems
We want now to extend the analysis to the multi-class case, i.e. to problems with $N_{C}>2$. Following the framework we introduced earlier, we can easily extend the computation. We will again consider the distribution among the $N_{C}$ classes of the dataset elements after initialization. In particular, we can define $N_{C}$ values associated with the fraction of dataset elements classified as belonging to each of the $N_{C}$ classes, $\left\{G_{i}\right\}_{i=0}^{N_{C}-1}$. We could also consider the distribution of the sorted frequencies; in other words, in each experiment, we order the classes according to the corresponding frequencies and not by their label. To distinguish these frequencies (and their statistics) from the set $\left\{G_{i}\right\}_{i=0}^{N_{C}-1}$ , we use indices $\tilde{i}$, and call this new set $\left\{G_{\tilde{i}}\right\}_{\tilde{i}=0}^{N_{C}-1}$. In particular $G_{\tilde{0}}$ indicates the biggest frequency among the $N_{C}$, $G_{\tilde{1}}$ the second one and so on. Finally we add the set (or subset) cardinality, $M$, on the output nodes set $\left\{O_{c,M}\right\}_{c=0}^{M-1}$. [^19] Here, the cardinality of the set indicates the number of classes over which the statistics are computed. The statistics of the r.v. change with the number of classes (or nodes), i.e., with the cardinality. Analogously to the set of fractions, we define the set of ranked output nodes $\left\{O_{\tilde{c},M}\right\}_{\tilde{c}=0}^{M-1}$ In the following, similarly to Sec. E.1 we will derive the distribution of $G_{0}$.
Figure H.1: The distribution $f_{G_{0}}\left({g_{0}}\right)$ in a multi-class problem with $N_{C}=10$. The empirical mean value is reported; note that despite the difference in the distribution the mean value is the same that we would observe using a linear activation function $\mkern 1.5mu\overline{\mkern-1.5muG_{0}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{N_{C}}$. For these simulations we used the dataset GB on the model SHLP with max pooling ($k=20$). (see App. I.4 for more details).
We will make in the derivation of $G_{0}$ the following approximation: We will repeat the same approach of the binary setting comparing the generic class ”0” with the class with bigger mean output value among the $N_{C}-1$ remaining classes.
The first ingredient that we need is therefore the statistics of $O_{\tilde{0},N_{C}-1}$. From the analysis of previous sections, we know that $f_{O_{\tilde{0},N_{C}-1}}^{(\chi)}\left({o}\right)$ is asymptotically Gaussian. The mean value of this Gaussian variable will follow the distribution of the maximum among $N_{C}-1$ drawn from the Gaussian distribution $f_{\mu_{c}}\left({m}\right)$ as, by definition,
$\left\langle{O_{\tilde{0},N_{C}-1}}\right\rangle\equiv\max_{c\in\{1,\dots,N_{C}\}}\left\langle{O_{c}}\right\rangle\,.$ (211)
@@
$f_{Y}\left({y}\right)\equiv f_{\left\langle{O_{\tilde{0},N_{C}-1}}\right\rangle}\left({y}\right)=\left(N_{C}-1\right)\mathcal{N}\left(y;0,\tilde{\sigma}^{2}\right)\left(\frac{1}{2}\left(1+\operatorname{erf}\left(\frac{y}{\sqrt{2}\tilde{\sigma}}\right)\right)\right)^{N_{C}-2}\,.$ (214)
@@
$f_{G_{0}}\left({g_{0}}\right)=\int_{-\infty}^{\infty}f_{\left\langle{O_{0,N_{C}}}\right\rangle}\left({\tilde{x}}\right)f_{\left\langle{O_{\tilde{0},N_{C}-1}}\right\rangle}\left({\tilde{x}-\hat{\Delta}\left(g_{0}\right)}\right)d\tilde{x}\,$, (215)
where $\left\langle{O_{\tilde{0},N_{C}-1}}\right\rangle$ the biggest mean output among the $N_{C}-1$. Note that $\hat{\Delta}\left(g_{0}\right)$ is a function of $g_{0}$. Substituting Eq. (212) and Eq. (214) we have:
$f_{G_{0}}\left({g_{0}}\right)=\int_{-\infty}^{\infty}\mathcal{N}\left(\tilde{x};0,\hat{\sigma^{2}}\right)(N_{C}-1)\mathcal{N}\left(\tilde{x};\hat{\Delta}\left(g_{0}\right),\hat{\sigma^{2}}\right)\left(\frac{1}{2}\left(1+\operatorname{erf}\left(\frac{\tilde{x}-\hat{\Delta}\left(g_{0}\right)}{\sqrt{2}\hat{\sigma}}\right)\right)\right)^{N_{C}-2}d\tilde{x}\,$. (216)
@@ H.1 Increasing number of classes exacerbates IGB
Having a large number of classes increases the probability that the center of the distribution related to one of these classes is an outlier of the distribution $f_{\mu_{c}}\left({m}\right)(x)$. However, since $f_{\mu_{c}}\left({m}\right)$ is Gaussian (a fast-decaying distribution), the typical value of the maximum increases slowly. Therefore, an extremely large number of classes, $N_{c}$, will be needed to observe significant changes.
We can analyze this quantitatively. The typical value of the maximum out of $N_{c}$ Gaussian random variables of variance $\epsilon$ grows as $\sqrt{(2\epsilon\log(N_{c}))}$ [Gumbel, 1958, Hartarsky et al., 2019]. In the absence of IGB, $\epsilon$ is the variance of $f_{\mu_{c}}\left({m}\right)$ and scales inversely with the dataset size, i.e., $\sim 1/D$ (see e.g., App. D.2 and Fig. D.1). This implies that outliers will only appear if the number of classes far exceeds the number of datapoints, a situation that does not occur in real life.
@@ I.1.1 A prototype for highly correlated data: MNIST
Figure I.1: Comparison of the distribution $f_{G_{0}}\left({g_{0}}\right)$ between two untrained neural networks fed with MNIST dataset; the ten classes of the dataset were merged in two macro groups: even and odd. The two architectures employed in the comparison differ in the choice of activation function and pooling. For these simulations we used the E&O dataset on the model CNN-B. (see App. I.4 for more details). On the right, the same plot in a logarithmic scale to better visualize the differences in the low-density regions.
@@ I.1.2 A further example: CIFAR10
Figure I.2: Comparison of the distribution $f_{G_{0}}\left({g_{0}}\right)$ between two untrained neural networks fed with CIFAR10 dataset (with 10 classes). The two architectures employed in the comparison differ in the choice of activation function and pooling. In the absence of IGB, the distributions would concentrate around the vertical line. For these simulations we used the dataset C10 on the model CNN-B described below (see App. I.4). On the right, the same plot in a logarithmic scale to better visualize the differences in the low-density regions.
We now show a multi-class example, on the same network as in App. I.1.1 (CNN-B) we propagated CIFAR10. The results are reported in Fig. I.2. As with MNIST, the results are qualitatively consistent with the conclusions of our theory: the network with ReLU and max pooling displays a stronger IGB than its counterpart with $\tanh$ and AvgPooling. Note, also, how in both cases, symmetry between classes is preserved at the ensemble level. In both cases, in fact, we get $\mkern 1.5mu\overline{\mkern-1.5muG_{c}\mkern-1.5mu}\mkern 1.5mu=1/N_{C}$ . To show this we show in Fig. I.2 two vertical lines at the mean values calculated on the empirical distributions (histograms), each surrounded by its own uncertainty, estimated through the standard error.
@@ I.1.3 High cardinality dataset: Cifar100
We also provide another example of a multi-class dataset. In this case, we choose Cifar-100 as a prototype of a high-cardinality dataset (with a high number of classes) to demonstrate the presence of IGB in this scenario as well. In the absence of IGB, the distribution should be peaked at $G_{0}=1/N_{C}=1/100$. However, the histogram shows a peak at 0 (most of the time, the generic class does not receive any assigned elements). The gap between the peak at 0 and the rest of the distribution, particularly evident in the log-log scale plot, indicates a resolution limit due to the finite number of dataset elements ($D=10^{4}$).
Figure I.3: CNN-A (ReLU + max pool case) with C100 in this case ($N_{C}=100$). The peak at $G_{0}=\frac{1}{N_{C}}$, corresponding to the No IGB case (in the absence of IGB, the distribution should concentrate around this value), is reported as a reference. The plot on the right contains the same data as that on the left, but has a logarithmic $x$ axis. We added a small numer, $\epsilon=1e-9$, to the value of $G_{0}$ in order to show the peak at $G_{0}=0$.
Figure I.4: Right: Comparison of the trend of $\max_{i}\{G_{i}\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). The curves show a consistent pattern with the diagram on the shown in Fig. 2. The simulations were conducted on a ResNet (top) and Vision Transformer (bottom) using a binary dataset (dogs vs cats from CIFAR) as input; more details on the setting are provided in App. I.4.
Figure I.5: Right: Comparison of the trend of $\max_{i}\{G_{i}\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). The curves show a consistent pattern with the diagram on the shown in Fig. 2. The simulations were conducted on different architectures (MLP- mixer (top), ResNet (middle) and Vision Transformer (bottom)) using a multi-class dataset (C10) as input; more details on the setting are provided in App. I.4.
@@ I.1.4 Additional architectures
We also provide empirical evidence of IGB in different architectures. To improve the legibility of the paper, we show results for ResNet, ViT and MLP-mix in Sec. I.2. While App. I.2 is devoted to the dynamics und IGB, the reader can recognize the presence of IGB in Figs. I.4 and I.5. Specifically, IGB is identified by checking the quantity $\max_{c}\{G_{c}(\mathcal{W}_{t})\}$, for $t=0$.; we can observe that at time $t=0$, i.e., at initialization, each of these curves shows $\max_{c}\{G_{c}\}\neq 1/N_{C}$. [^21] The reader might have particular interest in the $K=0$ case, where the data is centered and IGB cannot be attributed to data preprocessing.
@@ I.2 Experiments on other architectures & effects on the training dynamics
Even in the case of $K=0$, the system exhibits IGB, as indicated by $\max_{i}\{G_{i}\}\neq 1/N_{C}$ at initialization. Also, it is worth noticing that the susceptibility of networks to IGB varies, as seen from the comparison of the $\gamma(\mathcal{A},\psi(\chi))$ range in the colormaps of different experiments. Despite using the same set of $K$ values, the resulting IGB level varies significantly across networks. The IGB values, dependent on the combined effect of $\psi(\chi)$ and $\mathcal{A}$, show substantial variation among the considered architectures, suggesting the presence of regulatory or amplifying elements for IGB within these architectures.
@@ I.3 IGB in Pre-Trained Models
Figure I.6: The distribution $f_{G_{0}}\left({g_{0}}\right)$ on pre-trained models. From left to right are shown PT-ResNet, PT-SwinT-FC, PT-ENetV2-FC on a binary dataset (Cats and Dogs from C10).
@@
Figure I.7: The distribution $f_{G_{0}}\left({g_{0}}\right)$ on pre-trained models. From left to right are shown PT-SwinT-MLP, PT-ENetV2-MLP on a binary dataset (Cats and Dogs from C10)